# Negotiative Alignment: Embracing Disagreement to Achieve Fairer Outcomes – Insights from Urban Studies

### v1

@@ 4.1. Overview of Street-Level Findings
Figure 2. Matrix of compiled ratings for 20 Montreal streets across multiple dimensions (e.g., inclusivity, accessibility, aesthetics, practicality). Ratings are disaggregated by core demographic groups, illustrating distinct disagreements.
[…]
Matrix of compiled ratings for 20 Montreal streets
@@
4.2. Who Disagrees and Why
4.2.1. Wheelchair User Discrepancies
@@
4.2.2. LGBTQIA2+ Aesthetic Divergence
@@
4.2.4. Post-occupancy vs. Pre-occupancy
@@
Figure 3. Violin Plot of Inclusivity Ratings: Distribution of inclusivity ratings for 20 streets across demographic groups, highlighting differences in perceived inclusivity and the influence of features such as seating and sidewalk conditions.
[…]
Violin Plot of Inclusivity Ratings
4.3. Negotiation in Group Settings
@@
4.4. Rating vs. Ranking: Contrasting Insights
@@
  • Rating. The 4-point scale allowed for partial convergence on less contested dimensions but often masked deeper disagreements. This finding aligns with social choice theory, where rating methods capture the intensity of preferences but may overlook nuanced conflicts (Leroy et al., 2024; Xu et al., 2024; Rofin et al., 2023).
@@
  • Ranking. Forced-order exercises exposed sharp conflicts when participants were required to assign a single linear priority. For instance, disagreements between accessibility and aesthetics became more pronounced. This observation corroborates previous studies that indicate ranking methods can reveal underlying conflicts more effectively than rating methods (Suzuki and Horita, 2024; Fritz et al., 2023).
@@
Although negotiation could clarify some misunderstandings, the data reveal that certain trade-offs persist. This pattern supports an approach that retains disagreements in computational models instead of converging on an artificial consensus, consistent with recommendations in the social choice literature (Fritz et al., 2023).
4.5. Summary of Key Observations
  1. (1)
@@ 4.5. Summary of Key Observations
    Disagreement is Structurally Patterned. Divergences in urban evaluations are not random but systematically reflect structural inequities, particularly for marginalized groups (e.g., mobility, symbolic safety) (Zhu et al., 2025; Adams and Laursen, 2007; Cruz et al., 2018; Mela, 2014; Leroy et al., 2024; Jackson et al., 2016; Mushkani and Ono, 2021). These findings align with intersectional urban studies, demonstrating that public space assessments are shaped by overlapping exclusions based on race, gender, disability, and socioeconomic status. Our study extends this literature by formalizing these divergences within a computational framework for AI alignment.
  2. (2)
@@
    Ratings and Rankings Capture Distinct Information. Ratings offer finer granularity, allowing for the representation of intersectional positionalities, whereas rankings impose explicit trade-offs, often simplifying multi-dimensional concerns (Xu et al., 2024; Rofin et al., 2023). While prior research has conceptually distinguished these methods, our findings provide empirical validation of their respective strengths and limitations, highlighting the need for hybrid approaches in preference alignment.
  3. (3)
@@
    Negotiation Partially Mitigates Discrepancies. While no single urban site received unanimous evaluation, negotiation facilitated a more nuanced and balanced assessment. As shown in the violin plot (3), negotiative evaluations of urban inclusivity exhibit calibration effects at both individual and group levels. Objective clarifications help refine understanding, yet fundamental value-based disagreements persist across demographic lines, reinforcing the notion that negotiation alone cannot fully resolve underlying value conflicts (Suzuki and Horita, 2024). This dynamic reflects the inherent nature of cities as arenas of continuous negotiation among diverse individuals and entities (Lefebvre, 1974; Jacobs, 1961; Harvey, 1989).
  4. (4)
@@
    Averages Overlook Marginalization. A mean rating can obscure experiences of groups who find a space unusable or exclusionary, reinforcing concerns about aggregative methods masking minority viewpoints (Fritz et al., 2023).
@@ 4.6. Further Analyses
Figure 4. Rating using Pearson correlation coefficients between demographic groups for practicality, inclusivity, aesthetics, and accessibility. Higher disagreement is observed in accessibility and practicality scores, particularly among participants with intersectional identities. Notably, the individual identifying as female, Black, senior, and disabled, as well as the individual identifying as female, minority, adult, and disabled, exhibit greater divergence from the rest.
[…]
Rating using Pearson correlation
Figure 5. Ranking using Kendall Tau coefficients for accessibility, invitingness, comfort, and regenerativeness. Certain criteria, such as regenerativeness, show higher consensus, whereas accessibility remains more contentious.
[…]
Ranking using Kendall Tau3
Figure 6. Ranking using Kendall Tau for beauty, practicality, maintenance, and inclusivity. Disagreement spikes in practicality and inclusivity, reflecting differences in lived experiences.
[…]
Ranking using Kendall Tau2
Figure 7. Ranking using Kendall Tau for dynamism, representativeness, oppressiveness, and security. These dimensions often polarize participants, reflecting divergent perceptions of safety, vibrancy, and cultural representation.
[…]
Ranking using Kendall Tau
Figure 8. Group Rating using Pearson correlation coefficients for selected citizen groups (F1G, F2G, F3G) across practicality, inclusivity, aesthetics, and accessibility. The results highlight a more balanced evaluation.
[…]
Group Rating
5. Negotiative Alignment Framework
Our findings corroborate existing social choice theory regarding the strengths and limitations of rating and ranking methods (Leroy et al., 2024; Xu et al., 2024; Rofin et al., 2023; Suzuki and Horita, 2024; Fritz et al., 2023). Specifically, ratings captured the intensity of preferences, aligning with theories that emphasize their ability to reflect nuanced stakeholder values, while rankings effectively exposed fundamental conflicts, as previously documented. Beyond these confirmations, our study extends the literature by demonstrating how minorities with intersectional identities shape the dynamics of agreement and disagreement. For instance, wheelchair users and LGBTQIA2+ participants exhibited distinct preferences that traditional aggregation methods might overlook (Leroy et al., 2024; Xu et al., 2024). This highlights the necessity of incorporating negotiative alignment into AI models, which treats these persistent disagreements as meaningful inputs rather than statistical noise. By doing so, AI systems can better preserve minority viewpoints and adapt to the complex, evolving needs of diverse urban populations (Zhou et al., 2024; Kirk et al., 2024; Sorensen et al., 2024; Mushkani et al., 2025b).
5.1. Multi-Agent Utility and Constraint Formulation
@@
subject to $\sum\lambda_{g}=1,\,\lambda_{g}\geq 0$. This averaging tends to underrepresent minority experiences (Rawls, 2001).
5.1.1. Dynamic Multi-Agent Negotiation
@@
where $\Delta_{g}(t)$ captures newly surfaced disagreements at iteration $t$, and $C$ includes resource constraints or policy requirements (Moulin, 2004). The system re-optimizes after each iteration, reflecting evolving stakeholder concerns.
5.1.2. Iterative Negotiation Operator
@@
which identifies critical disagreements, proposes possible design or weighting adjustments, and employs a rule such as Nash or Kalai-Smorodinsky bargaining (Nash, 1950; Kalai and Smorodinsky, 1975) to guide solutions.
5.2. Identity Preservation and Fairness Metrics
@@
where $\hat{\pi}_{g}$ is the updated preference distribution, and $\beta$ controls how quickly preferences adapt (Zinkevich, 2003).
@@
5.3. Benchmarks for Evaluating Negotiative Alignment
@@
5.4. Prototype Demonstration
To illustrate negotiative alignment, we implemented a small-scale prototype on subset Vivian Avenue (i09) (inclusivity dimension). Table 1 shows each group’s rating, and Table 3 presents the iterative consensus after two negotiation rounds. Our Python code (see Appendix A.6) updates group weights using disagreement magnitudes while preserving identity distributions via a Jensen–Shannon-based metric.
@@ 5.4. Prototype Demonstration
Table 1. Stakeholder ratings for urban scene i09.
@@
Table 2. Weight adjustments across iterations in the negotiation process.
@@
Table 3. Consensus value $x^{*}$, Disagreement Coverage Ratio (DCR), and Identity Preservation Index (IPI) during negotiation iterations.
Participants’ group rating for i09 yielded a score of 2.6. Compared to a naive average, the negotiative alignment is a small step closer to that group rating. These results show that our framework can refine consensus while preserving key aspects of each group’s original preferences. Further studies and testing are required.
5.5. Implications for Fairness and Accountability
@@
5.6. Philosophical and Ethical Considerations
Negotiative alignment views persistent value disagreements as legitimate, rather than anomalies to be averaged out. This resonates with pluralistic conceptions of justice and fairness (Sen, 2009). However, practical deployment requires careful attention to how negotiation power is distributed among stakeholders. Mechanism design strategies may help allocate bargaining power more equitably, ensuring no single group unilaterally dictates outcomes.
5.7. Limitations and Future Directions
@@ Appendix A Supplementary Material
A.1. Heatmap of Correlations by Demographic Group
Figure 9. Correlation heatmap illustrating the relationships between ratings for 20 urban sites across four evaluation criteria—practicality, inclusivity, aesthetics, and accessibility—segmented by demographic groups (LGBTQIA2+, individuals with disabilities, elderly females, elderly males, young males, young females, and group-based evaluations). We observe that practicality and accessibility tend to exhibit stronger correlations within most groups, suggesting some alignment in how these dimensions are evaluated. In contrast, correlations involving aesthetics and inclusivity are generally weaker or negative, possibly reflecting divergent perceptions or trade-offs between these criteria. Group-based evaluations show greater consistency with individual ratings for inclusivity and practicality compared to aesthetics, which appears more variable. Notable differences between demographic groups, such as weaker correlations between accessibility and inclusivity for young males and individuals with disabilities, may indicate varying priorities or lived experiences. These findings suggest the need for AI models to account for diverse perspectives when interpreting such evaluation data.
[…]
Correlation heatmap
A.2. Criteria Rating Table
Figure 10. Criteria Rating Table: Summary of the four evaluation criteria and their corresponding rating levels. A score of 1 represents the lowest level of performance, while a score of 4 represents the highest. Participants used a sticker-based rating system with color coding to visually represent their assessments. This method allowed for straightforward comparison of ratings across the selected criteria.
[…]
@@ A.2. Criteria Rating Table
A.3. Representative Street Views
Figure 11. Representative Street Views: A selection of street-level images from 20 Montreal neighborhoods, capturing variations in land use, pedestrian infrastructure, greenery, and streetscape design. These images provide visual context for the urban spaces included in the study and help to illustrate the diversity of the evaluated sites.
[…]
@@ A.3. Representative Street Views
A.4. Spatial Distribution of Study Sites
Figure 12. Spatial Distribution of Study Sites: Map of the geographic locations of the 20 selected urban sites across Montreal. We selected these sites to represent a broad spectrum of socio-economic contexts, land uses, urban densities, and historical characteristics. The spatial distribution highlights the diversity of the sites included in our analysis. The base map is sourced from OpenStreetMap.
[…]
@@ A.4. Spatial Distribution of Study Sites
A.5. Street Diversity Classification Guide
Figure 13. Street Diversity Classification Guide: Matrix categorizing the selected streets based on socio-spatial attributes, including land use type, history, greenery, and aesthetic features. Our selection process aimed to ensure diversity by incorporating at least one representative street for each category. For example, streets were chosen to include residential, commercial, and mixed-use areas, as well as variations in history and vegetation levels. This approach provided a structured yet flexible framework for selecting sites that reflect the heterogeneity of Montreal’s urban environment.
[…]
A matrix categorizing streets
A.6. Prototype Demonstration Code

### v2

@@ Results
Figure 2: Matrix of compiled ratings for 20 Montreal streets across multiple dimensions (e.g., inclusivity, accessibility, aesthetics, practicality). Ratings are disaggregated by core demographic groups (each group consists of 2 persons), illustrating distinct disagreements.
@@
Who Disagrees and Why
Wheelchair User Discrepancies
@@
LGBTQIA2+ Aesthetic Divergence
@@
Post-occupancy vs. Pre-occupancy
@@
Figure 3: Violin Plot of Inclusivity Ratings: Distribution of inclusivity ratings for 20 streets across demographic groups, highlighting differences in perceived inclusivity and the influence of features such as seating and sidewalk conditions.
Negotiation in Group Settings
@@
Rating vs. Ranking: Contrasting Insights
@@
  • Rating. The 4-point scale allowed for partial convergence on less contested dimensions but often masked deeper disagreements. This finding aligns with social choice theory, where rating methods capture the intensity of preferences but may overlook nuanced conflicts (Leroy et al. 2024; Xu et al. 2024; Rofin et al. 2023).
@@
  • Ranking. Forced-order exercises exposed sharp conflicts when participants were required to assign a single linear priority. For instance, disagreements between accessibility and aesthetics became more pronounced. This observation corroborates previous studies that indicate ranking methods can reveal underlying conflicts more effectively than rating methods (Suzuki and Horita 2024; Fritz, Moretti, and Staudacher 2023).
@@
Although negotiation could clarify some misunderstandings, the data reveal that certain trade-offs persist. This pattern supports an approach that retains disagreements in computational models instead of converging on an artificial consensus, consistent with recommendations in the social choice literature (Fritz, Moretti, and Staudacher 2023).
Summary of Key Observations
  1. 1.
@@
    Disagreement is Structurally Patterned. Divergences in urban evaluations are not random but systematically reflect structural inequities, particularly for marginalized groups (e.g., mobility, symbolic safety) (Zhu, Zhang, and Biljecki 2025; Adams and Laursen 2007; Cruz, Roskamm, and Charalambous 2018; Mela 2014; Leroy et al. 2024; Jackson, Williams, and VanderWeele 2016; Mushkani and Ono 2021). These findings align with intersectional urban studies, demonstrating that public space assessments are shaped by overlapping exclusions based on race, gender, disability, and socioeconomic status. Our study extends this literature by formalizing these divergences within a computational framework for AI alignment.
  2. 2.
@@
    Ratings and Rankings Capture Distinct Information. Ratings offer finer granularity, allowing for the representation of intersectional positionalities, whereas rankings impose explicit trade-offs, often simplifying multi-dimensional concerns (Xu et al. 2024; Rofin et al. 2023). While prior research has conceptually distinguished these methods, our findings provide empirical validation of their respective strengths and limitations, highlighting the need for hybrid approaches in preference alignment.
  3. 3.
@@
    Negotiation Partially Mitigates Discrepancies. While no single urban site received unanimous evaluation, negotiation facilitated a more nuanced and balanced assessment. As shown in the violin plot (3), negotiative evaluations of urban inclusivity exhibit calibration effects at both individual and group levels. Objective clarifications help refine understanding, yet fundamental value-based disagreements persist across demographic lines, reinforcing the notion that negotiation alone cannot fully resolve underlying value conflicts (Suzuki and Horita 2024). This dynamic reflects the inherent nature of cities as arenas of continuous negotiation among diverse individuals and entities (Lefebvre 1974; Jacobs 1961; Harvey 1989).
  4. 4.
@@
    Averages Overlook Marginalization. A mean rating can obscure experiences of groups who find a space unusable or exclusionary, reinforcing concerns about aggregative methods masking minority viewpoints (Fritz, Moretti, and Staudacher 2023).
@@
Figure 4: Rating using Pearson correlation coefficients between demographic groups for practicality, inclusivity, aesthetics, and accessibility. Higher disagreement is observed in accessibility and practicality scores, particularly among participants with intersectional identities. Notably, the individual identifying as female, Black, senior, and disabled, as well as the individual identifying as female, minority, adult, and disabled, exhibit greater divergence from the rest.
Figure 5: Ranking using Kendall Tau coefficients for accessibility, invitingness, comfort, and regenerativeness. Certain criteria, such as regenerativeness, show higher consensus, whereas accessibility remains more contentious.
Figure 6: Ranking using Kendall Tau for beauty, practicality, maintenance, and inclusivity. Disagreement spikes in practicality and inclusivity, reflecting differences in lived experiences.
Figure 7: Ranking using Kendall Tau for dynamism, representativeness, oppressiveness, and security. These dimensions often polarize participants, reflecting divergent perceptions of safety, vibrancy, and cultural representation.
Figure 8: Group Rating using Pearson correlation coefficients for selected citizen groups (F1G, F2G, F3G) across practicality, inclusivity, aesthetics, and accessibility. The results highlight a more balanced evaluation.
Negotiative Alignment Framework
Our findings corroborate existing social choice theory regarding the strengths and limitations of rating and ranking methods (Leroy et al. 2024; Xu et al. 2024; Rofin et al. 2023; Suzuki and Horita 2024; Fritz, Moretti, and Staudacher 2023). Specifically, ratings captured the intensity of preferences, aligning with theories that emphasize their ability to reflect nuanced stakeholder values, while rankings effectively exposed fundamental conflicts, as previously documented. Beyond these confirmations, our study extends the literature by demonstrating how minorities with intersectional identities shape the dynamics of agreement and disagreement. For instance, wheelchair users and LGBTQIA2+ participants exhibited distinct preferences that traditional aggregation methods might overlook (Leroy et al. 2024; Xu et al. 2024). This highlights the necessity of incorporating negotiative alignment into AI models, which treats these persistent disagreements as meaningful inputs rather than statistical noise. By doing so, AI systems can better preserve minority viewpoints and adapt to the complex, evolving needs of diverse urban populations (Zhou et al. 2024; Kirk et al. 2024; Sorensen et al. 2024; Mushkani et al. 2025b).
Multi-Agent Utility and Constraint Formulation
@@
subject to $\sum\lambda_{g}=1,\,\lambda_{g}\geq 0$. This averaging tends to underrepresent minority experiences (Rawls 2001).
Dynamic Multi-Agent Negotiation
@@
where $\Delta_{g}(t)$ captures newly surfaced disagreements at iteration $t$, and $C$ includes resource constraints or policy requirements (Moulin 2004). The system re-optimizes after each iteration, reflecting evolving stakeholder concerns.
Iterative Negotiation Operator
@@
which identifies critical disagreements, proposes possible design or weighting adjustments, and employs a rule such as Nash or Kalai-Smorodinsky bargaining (Nash 1950; Kalai and Smorodinsky 1975) to guide solutions.
Identity Preservation and Fairness Metrics
@@
where $\hat{\pi}_{g}$ is the updated preference distribution, and $\beta$ controls how quickly preferences adapt (Zinkevich 2003).
@@
Benchmarks for Evaluating Negotiative Alignment
@@
Prototype Demonstration
To illustrate negotiative alignment, we implemented a small-scale prototype on subset Vivian Avenue (i09) (inclusivity dimension). Table 1 shows each group’s rating, and Table 3 presents the iterative consensus after two negotiation rounds. This serves as a microcosm of the full analysis across 20 urban scenes (Appendix A7). Our Python code (Appendix A6) updates group weights based on disagreement magnitudes while preserving identity distributions via a Jensen–Shannon-based metric.
@@
Table 1: Stakeholder ratings for urban scene i09.
@@
Table 2: Weight adjustments across iterations in the negotiation process.
@@
Table 3: Consensus value $x^{*}$, Disagreement Coverage Ratio (DCR), and Identity Preservation Index (IPI) during negotiation iterations.
During the empirical study, participants’ group rating for i09 was 2.6. The naive average, computed across all participants, was 2.067, while the negotiative alignment score was 2.224, reflecting a moderate shift toward the group rating. Detailed calculations for all 20 urban scenes—evaluated for inclusivity, accessibility, practicality, and aesthetics—are provided in Appendix A7. These findings illustrate how selecting appropriate hyperparameters and iteration settings can affect the resulting negotiated score. The robustness of the final output depends on both the quantity and specificity of agent-related data, as well as on parameter tuning within the model. This iterative weighting process further clarifies the relative emphasis placed on each group’s viewpoint.
This prototype demonstrates the framework’s capacity to refine consensus while preserving critical elements of each group’s initial preferences. However, it relies on a single criterion and participant demographic data, without additional contextual or agent-specific information. Further evaluation is necessary to verify the framework’s applicability under more comprehensive conditions.
Integrating Large Language Models (LLMs) into the Negotiative Alignment Loop
The negotiative alignment framework can be instantiated with LLMs by treating each stakeholder group as an LLM agent, whose preferences are encoded through prompts or fine-tuning. Each agent receives a summary of the current design proposal and a concise representation of other agents’ positions, then generates a counter-proposal or weight adjustment. Because LLMs have finite context windows, the negotiation protocol must specify which aspects of the design and which past exchanges are carried forward. In practice, this can be achieved by maintaining a rolling buffer of key arguments or by distilling earlier rounds into abstracted summaries that preserve points of contention without exceeding the model’s input limits.
Negotiation proceeds either in an iterative manner—where agents exchange multiple rounds of proposals to approximate middle grounds—or as a one-shot exchange—where all relevant information is compressed into a single prompt. An iterative protocol can mitigate context constraints by revealing information gradually, thereby allowing the model to focus on the most salient disagreements at each step. Conversely, a one-shot protocol may be preferable when the number of stakeholders is small and the full set of arguments fits within the LLM’s input capacity. In both cases, the negotiation operator $\mathcal{N}$ is realized as a sequence of prompt templates that apply formal bargaining rules (e.g., Nash or Kalai–Smorodinsky) to the agents’ generated weight adjustments.
Within each round, LLM agents can apply internal voting mechanisms to generate collective proposals. A majority-voting prompt instructs the model to select the most supported design features, while preference-ranking prompts ask the model to order options by stakeholder-weighted utility. These procedures can be embedded directly into the prompt engineering, enabling the LLM to simulate both deliberation and decision aggregation without external algorithms. The resulting proposals and weight updates are then fed back into the negotiation loop, closing the chain of prompts that implements the iterative operator.
The outputs of each LLM-mediated round also yield text-based records that facilitate the computation of fairness metrics. For example, identity preservation can be measured by comparing the distribution of model-reported preferences at each step against the original set of stakeholder prompts, using divergence measures such as Jensen–Shannon. The degree of negotiation, or the extent of weight adjustment between rounds, can be quantified by tracking the absolute differences in the weight allocations reported by the LLM. By leveraging the generative and reasoning capabilities of large language models, the framework supports a transparent, auditable negotiation process in which disagreement is both preserved in the text transcripts and systematically addressed through successive prompt exchanges.
This paper has articulated the need for more effective negotiation methods in AI-driven decision-making and introduced negotiative alignment as a candidate framework. Future work should implement and empirically evaluate LLM-based negotiation protocols and design studies to assess negotiative alignment in practice. Further research is necessary to advance these methods across diverse application contexts.
Implications for Fairness and Accountability
@@
Philosophical and Ethical Considerations
Negotiative alignment views persistent value disagreements as legitimate, rather than anomalies to be averaged out. This resonates with pluralistic conceptions of justice and fairness (Sen 2009). However, practical deployment requires careful attention to how negotiation power is distributed among stakeholders. Mechanism design strategies may help allocate bargaining power more equitably, ensuring no single group unilaterally dictates outcomes.
Limitations and Future Directions
@@ Appendix A Supplementary Material
Heatmap of Correlations by Demographic Group
Figure 9: Correlation heatmap illustrating the relationships between ratings for 20 urban sites across four evaluation criteria—practicality, inclusivity, aesthetics, and accessibility—segmented by demographic groups (LGBTQIA2+, individuals with disabilities, elderly females, elderly males, young males, young females, and group-based evaluations). We observe that practicality and accessibility tend to exhibit stronger correlations within most groups, suggesting some alignment in how these dimensions are evaluated. In contrast, correlations involving aesthetics and inclusivity are generally weaker or negative, possibly reflecting divergent perceptions or trade-offs between these criteria. Group-based evaluations show greater consistency with individual ratings for inclusivity and practicality compared to aesthetics, which appears more variable. Notable differences between demographic groups, such as weaker correlations between accessibility and inclusivity for young males and individuals with disabilities, may indicate varying priorities or lived experiences. These findings suggest the need for AI models to account for diverse perspectives when interpreting such evaluation data.
@@
Figure 10: Criteria Rating Table: Summary of the four evaluation criteria and their corresponding rating levels. A score of 1 represents the lowest level of performance, while a score of 4 represents the highest. Participants used a sticker-based rating system with color coding to visually represent their assessments. This method allowed for straightforward comparison of ratings across the selected criteria.
@@
Figure 11: Representative Street Views: A selection of street-level images from 20 Montreal neighborhoods, capturing variations in land use, pedestrian infrastructure, greenery, and streetscape design. These images provide visual context for the urban spaces included in the study and help to illustrate the diversity of the evaluated sites.
@@
Figure 12: Spatial Distribution of Study Sites: Map of the geographic locations of the 20 selected urban sites across Montreal. We selected these sites to represent a broad spectrum of socio-economic contexts, land uses, urban densities, and historical characteristics. The spatial distribution highlights the diversity of the sites included in our analysis. The base map is sourced from OpenStreetMap.
Street Diversity Classification Guide
Figure 13: Street Diversity Classification Guide: Matrix categorizing the selected streets based on socio-spatial attributes, including land use type, history, greenery, and aesthetic features. Our selection process aimed to ensure diversity by incorporating at least one representative street for each category. For example, streets were chosen to include residential, commercial, and mixed-use areas, as well as variations in history and vegetation levels. This approach provided a structured yet flexible framework for selecting sites that reflect the heterogeneity of Montreal’s urban environment.
Prototype Demonstration Code